### v1

# Monocular 3D lane detection for Autonomous Driving: Recent Achievements, Challenges, and Outlooks

Abstract
3D lane detection plays a crucial role in autonomous driving by extracting structural and traffic information from the road in 3D space to assist the self-driving car in rational, safe, and comfortable path planning and motion control. Due to the consideration of sensor costs and the advantages of visual data in color information, in practical applications, 3D lane detection based on monocular vision is one of the important research directions in the field of autonomous driving, which has attracted more and more attention in both industry and academia. Unfortunately, recent progress in visual perception seems insufficient to develop completely reliable 3D lane detection algorithms, which also hinders the development of vision-based fully autonomous self-driving cars, i.e., achieving level 5 autonomous driving, driving like human-controlled cars. This is one of the conclusions drawn from this review paper: there is still a lot of room for improvement and significant improvements are still needed in the 3D lane detection algorithm for autonomous driving cars using visual sensors. Motivated by this, this review defines, analyzes, and reviews the current achievements in the field of 3D lane detection research, and the vast majority of the current progress relies heavily on computationally complex deep learning models. In addition, this review covers the 3D lane detection pipeline, investigates the performance of state-of-the-art algorithms, analyzes the time complexity of cutting-edge modeling choices, and highlights the main achievements and limitations of current research efforts. The survey also includes a comprehensive discussion of available 3D lane detection datasets and the challenges that researchers have faced but have not yet resolved. Finally, our work outlines future research directions and welcomes researchers and practitioners to enter this exciting field.
Figure 1: Chronological overview of the most prestigious monocular image-based 3D lane detection. For methods without abbreviations, we use circled numbers as substitutes, and the corresponding paper names are listed above the timeline chart.
I INTRODUCTION
In recent advancements fueled by artificial intelligence, autonomous driving technology has seen rapid evolution, progressively reshaping the paradigms of human transportation. Equipped with an array of sensors, autonomous vehicles emulate human sensory capabilities such as vision and hearing to perceive their surroundings and interpret traffic scenarios for safe navigation[98]. Key among these sensors are LiDAR, high-resolution cameras, millimeter-wave radars, and ultrasonic radars[119, 55], which facilitate feature extraction and object classification combined with high-precision mapping to discern obstacles and the vehicular traffic landscape[114]. Visual sensors, the most extensively utilized in autonomous vehicles, serve as the primary means for environmental perception, encompassing lane detection, traffic light analysis, road sign detection and recognition, vehicle tracking, pedestrian detection, and short-term traffic prediction[11]. The processing and understanding of visual scenes in autonomous driving, which includes analysis of traffic lights, identification of traffic signs, lane detection, and the detection of nearby pedestrians and vehicles, underpin more robust and safer commands for maneuvers such as steering, overtaking, lane changes, or braking[82, 56]. This integration of sensory data and environmental understanding seamlessly transitions into the domain of scene understanding in autonomous driving, which is pivotal for advancing vehicular autonomy and ensuring road safety.
Scene understanding represents one of the most challenging aspects within the domain of autonomous driving[57]. Without the capability for comprehensive scene understanding, navigating an autonomous vehicle safely through traffic lanes can be as daunting as navigating the world blindfolded for humans. Lane detection, in particular, stands out as a pivotal and challenging task within the realm of scene comprehension[105]. Lanes serve as the most common traffic elements on roads, acting as crucial markers that segment the roadways to ensure the safe and efficient passage of vehicles[79]. Lane detection technology, which automatically identifies road markings, is indispensable; autonomous vehicles lacking this capability could lead to traffic congestion and even severe collisions, thereby compromising passenger safety. Thus, lane detection plays a critical role in the ecosystem of autonomous driving. Unlike typical objects, lane markings occupy only a slender portion of the road scene and are widely distributed, making them uniquely challenging to detect[76]. The task is further complicated by a variety of lane markings, insufficient lighting, obstructions, and interference from similar textures, which are common in many driving scenarios, thus exacerbating the intrinsic challenges associated with lane detection.
Lane detection methods based on monocular vision can be primarily categorized into traditional manual feature approaches and deep learning-based methods[58]. Early endeavors primarily focused on extracting low-level manual features[28, 41], such as edges and color information. However, these approaches often involve complex feature extraction and post-processing designs, and exhibit limited robustness in dynamically changing scenes. Traditional lane detection algorithms based on manual feature extraction commence by identifying characteristics such as color, texture, edges, orientation, and shape of the lane lines, followed by constructing detection models that approximate the lane markings as straight or higher-order curved lines[117]. Nevertheless, due to the lack of distinguishing features and poor adaptability to dynamic environments, traditional methods based on manual features are typically less reliable and more computationally expensive[96]. With the rapid evolution of deep learning, significant strides have been made in the fields of image classification, object detection, and semantic segmentation within computer vision, bringing innovative perspectives to the research of lane detection[97]. Deep neural networks (DNNs), rooted in deep learning, demonstrate profound capability in feature extraction from image data, with Convolutional Neural Networks (CNNs) being the most extensively applied. CNNs represent a specialized category of DNNs, characterized by multiple convolutional layers and bases, making them particularly suitable for processing structured data like visual images, and providing efficient feature extraction for various subsequent tasks[106]. In the context of lane detection, this translates to utilizing deep CNNs to extract high-level features in real-time frames, which are then processed by the model to accurately determine the positions of lane lines[31].
I-A Background and Related Works
Due to advancements in deep learning technologies, researchers have developed numerous strategies that have significantly simplified, expedited, and enhanced the task of lane detection[78]. Concurrently, as deep learning becomes increasingly widespread and as new concepts continuously emerge[59], the domain of lane detection has seen a further specialization and refinement in its approaches. Reflecting on the mainstream research trajectories within this domain, camera-based lane detection methods can be principally segmented into two-dimensional (2D) and three-dimensional (3D) lane detection paradigms[12].
2D lane detection. [35, 51] methods aim to accurately delineate the shape and position of lanes within images, primarily employing four distinct approaches: segmentation-based, anchor-based, keypoint-based, and curve-based strategies. (1) Segmentation-based approaches frame 2D lane detection as a pixel-level classification challenge, generating lane masks[116, 64, 65, 116]. These methods cultivate lane groupings through the exploration of effective semantic features and subsequent post-processing, although they are marked by their high computational costs. (2) Anchor-based approaches are lauded in 2D lane detection for their simplicity and efficiency, typically utilizing linear anchor points to regress positional offsets relative to targets [47, 78, 115]. To circumvent the constraints of linear anchors, [35]employs an intrinsic lane space generating various lane candidates. Heuristically designed row anchors classify row pixels as lanes [109, 50], further evolved into mixed (row and column) anchors in [37] to alleviate positioning errors for side lanes. Moreover, this method significantly enhances inference speed. (3) Keypoint-based approaches offer a more flexible and sparse modeling of lane positions[42, 67, 86], initiating with the estimation of point locations, followed by associating keypoints belonging to the same lane using varying schemes. The primary strategy in [72, 78] is to predict 2D lanes by representing lane structures with predefined keypoints and regressing offsets between sampled and predefined points. Despite promising results, these methods lack the flexibility to adapt to complex lane configurations due to their fixed point design. (4) Curve-based approaches [77, 83, 51]focus on fitting lane lines through various curve equations and specific parameters, transforming 2D lane detection into a curve parameter regression challenge by detecting start and end points alongside curve parameters. Although there have been promising advancements in 2D lane detection, a significant gap remains between the 2D outcomes and the real-world application requirements, notably the precise 3D positioning.
3D lane detection. Due to the inherent lack of depth information in 2D lane detection, projecting these detections into 3D space can result in inaccuracies and reduced robustness[8]. Consequently, numerous researchers have shifted their focus towards lane detection within the 3D domain. Deep learning-based 3D lane detection methodologies primarily bifurcate into CNN-based and Transformer-based approaches, initially constructing dense Bird’s Eye View (BEV) feature maps, followed by the extraction of 3D lane information from these intermediate representations[31].
CNN-based methods predominantly include D-LaneNet [24], which proposes a dual-path architecture utilizing Inverse Perspective Mapping (IPM) to transpose features and detect lanes through vertical anchor regression. 3D-LaneNet+ [16] segments BEV features into non-overlapping units, addressing the limitations of anchor orientation through lateral offsets, angles, and elevation shifts relative to the unit’s center. GenLaneNet [25] pioneers the use of a fictitious top-view coordinate system for better feature alignment and introduces a two-stage framework to decouple lane segmentation and geometric encoding. BEVLaneDet [91] employs a virtual camera to ensure spatial consistency and adapts to more complex scenarios through a keypoint-based representation of 3D lanes. GroupLane [48] innovates within the BEV by introducing a row-based classification method, which accommodates lanes in any direction and interacts with feature information within instance groups.
Transformer-based methods include CLGo [52], presenting a two-stage framework capable of estimating camera posture from the image and lane decoders based on BEV features. PersFormer [12] constructs dense BEV queries using offline camera poses, unifying 2D and 3D lane detection under a Transformer-based framework. STLanes3D [95] predicts 3D lanes using fused BEV features and introduces the 3DLane-IOU loss to couple lateral and height errors. Anchor3DLane [31], a CNN-based method, regresses 3D lanes directly from image features based on 3D anchors, significantly reducing computational overhead. CurveFormer [7] leverages sparse query representations and cross-attention mechanisms within Transformers to effectively regress polynomial coefficients of 3D lanes. LATR [53], building on CurveFormer’s query anchor modeling, constructs lane-aware query generators and dynamic 3D ground position embeddings. CurveFormer++[8] proposes a single-stage Transformer detection method that does not require image feature view transformation and directly infers 3D lane detection results from perspective image features.
I-B Challenges and Motivation
Accurately estimating the 3D position of lane markings demands robust depth perception, particularly in intricate urban settings with fluctuating lighting and weather conditions[43]. Furthermore, real-world data for 3D lane detection exhibits high variability due to factors such as diverse road types, markings, and environmental conditions, making it arduous to train models that generalize well across different scenarios[10]. Processing 3D data for lane detection entails substantial computational resources; this is especially critical in real-time applications where low latency is imperative. Additionally, lane markings may be obscured or degraded by various environmental factors such as occlusions, shadows, rain, or snow, presenting challenges for reliable detection in adverse conditions[113]. Moreover, integrating 3D lane detection into a comprehensive perception system alongside other sensors like cameras, LiDAR, and radar, and processing their combined outputs, poses integration challenges. Unfortunately,the community lacks a consolidated, unified, single point of reference for ascertaining the current level of maturity of camera-based 3D lane detection techniques for autonomous driving.
Considering the aforementioned challenges and the importance of vision sensor-based semantic segmentation in accurate scene understanding and parsing, we have accumulated existing research contributions and outcomes in this survey. The main research questions highlighted in this survey are as follows: 1) Do existing datasets possess the potential for 3D lane detection in complex visual scenes? 2) What are the model sizes and inference speeds of current methods, and can these methods meet the real-time requirements of autonomous vehicles? 3) Can current methods effectively perform 3D lane detection in complex visual scenes containing uncertainties such as fog and rain?
I-C Contributions
This survey takes a step ahead by critically examining the recent state-of-the-art in 3D lane detection techniques, and makes the following main contributions to the community:
  1. []
  2. 1.
    A comprehensive introduction to 3D lane detection techniques, defining the generic pipeline and explaining each step individually. This helps newcomers in the field to quickly grasp prior knowledge and contributions from previous research, especially in the context of autonomous driving. To the best of our knowledge, this is the first survey on camera-based 3D lane detection.
  3. 2.
    A discussion and critical analysis of the most relevant papers and datasets that have received significant attention in the field of 3D lane detection in recent years.
  4. 3.
    A performance study of current state-of-the-art methods, considering their computational resource requirements and the platforms for which these methods are developed.
  5. 4.
    A reasoned derivation of future research guidelines based on the analyzed literature, identifying open problems and challenges in this domain, as well as research opportunities that can be explored effectively to address them.
I-D Review Methodology
The research works discussed in this survey were retrieved using different keywords such as 3D lane detection in autonomous vehicles, vision-based 3D lane detection, and learning-based 3D lane detection. Most of the retrieved papers are directly relevant to the research topic, but there are also some exceptions, such as multi-modal methods [54, 6], and point cloud-based approaches [36], which are less related to the theme of this survey. Additionally, the aforementioned keywords were searched in multiple repositories, including Web of Science and Google Scholar, to ensure the retrieval of relevant content. Inclusion criteria ensured that a paper was recognized by experts in autonomous driving, based on factors such as citation count or the impact of prior work. It is worth mentioning that when consulting the literature, no work on monocular 3D lane detection based on traditional methods was found. This may be because, unlike single-camera 2D lane detection, which only requires identifying pixels belonging to the lane in a 2D image, monocular 3D lane detection requires determining the three-dimensional position information of the lane in the 3D space using the 2D image. Without the assistance of distance measuring sensors such as LiDAR  [^1] https://en.wikipedia.org/wiki/Lidar or prediction through deep learning, this is difficult to achieve.
The remainder of this paper is structured as follows: Section II provides an overview of related works for 3D lane detection in autonomous driving. Section III introduces performance evaluation of 3D lane detection. In section IV, commonly used datasets are introduced. And, Next, in section V, we introduce the current challenges and prospects for future 3D lane detection. Finally, Section VI concludes the paper.
Figure 2: An overview of the 3D-LaneNet architecture. Information is processed in two parallel streams or pathways: the image-view pathway and the top-view pathway. This is called the dual-pathway backbone.
II MONOCULDAR 3D LANE DETECTION IN AD
With the rapid development of deep learning and autonomous driving, deep learning-based monocular lane detection has received increasing attention from both the industry and academia. In the field of monocular lane detection, early work mainly focused on 2D lane detection. As autonomous driving technology matures, higher requirements are placed on lane detection, with the goal of predicting the three-dimensional information of lane lines from a single image. Therefore, starting from 2018, there have been successive works on monocular 3D lane detection. As shown in Fig. 1, the figure provides a chronological overview of the monocular 3D lane detection algorithms. It can be seen that with the passage of time, more and more works have emerged, indicating that this field is becoming increasingly popular. In this figure, the green arrows represent CNN-based methods, and the orange arrows represent Transformer-based methods.
Among these methods, 3D-LaneNet [24] is a pioneering work in the field of monocular 3D lane detection. 3D-LaneNet introduces a network that can directly predict three-dimensional lane information in road scenes from monocular images. This work is the first to solve the task of three-dimensional lane detection using on-board monocular vision sensor. 3D-LaneNet introduces two novel concepts: intra-network feature map inverse perspective mapping (IPM) and anchor-based lane representation. The intra-network IPM projection facilitates dual representation information flow in both front view and bird’s-eye view. The anchor-based lane output representation supports an end-to-end training approach, which is different from common heuristic methods that equate the problem of detecting three-dimensional lane lines to object detection problems. The overview of 3D-LaneNet is depicted in the Fig. 2. Inspired by works such as FCOS [81] and CenterNet [118], 3D LaneNet+ [16] is an anchor-free 3D lane detection algorithm that can detect 3D lane lines of arbitrary topological structures. The authors of 3D LaneNet+ followed the dual-stream network of 3D LaneNet that separately processes the image view and bird’s-eye view and extended it to support the detection of 3D lane lines with more topological structures. Instead of predicting the entire lane as a whole, 3D LaneNet+ detects small lane segments that lie within the cell and their attributes (position, orientation, height). In addition, the method learns a global embedding for each cell to cluster small lane segments into complete 3D lane information. In [33], the Jiang et al. designed a two-stage 3D lane detection network where each stage is trained separately. The first subnetwork focuses on lane image segmentation, while the second subnetwork focuses on predicting the 3D lane structure based on the segmentation outputs from the first subnetwork. The Efficient Channel Attention (ECA) [88] attention mechanism and the Convolutional Block Attention Module (CBAM) [101] attention mechanism were introduced in each stage, respectively, resulting in improved segmentation performance and accuracy in 3D lane detection. In [25], Guo et al. proposed Gen-LaneNet, a general and scalable 3D lane detection method for detecting 3D lane lines from a single image, as shown in Fig. 3. The authors introduce a novel geometry-guided lane anchor representation and apply specific geometric transformations to directly compute the real 3D lane points from the network output. This anchor design is an intuitive extension of the anchor design in 3D-LaneNet. The method aligns the anchor coordinates with the underlying bird’s-eye-view features, which makes it more robust in handling unfamiliar scenes. Additionally, the paper presents a scalable two-stage framework that allows independent learning of the image segmentation subnetwork and the geometry encoding subnetwork, significantly reducing the number of 3D labels required for training. Furthermore, the paper introduces a highly realistic synthetic image dataset with rich visual variations for the development and evaluation of 3D lane detection methods.
Figure 3: The Gen-LaneNet procedure involves encoding an input image into deep features using a segmentation backbone, which are then decoded into a lane segmentation map. The 3D-GeoNet then focuses on geometry encoding and predicts intermediate 3D lane points represented in top-view 2D coordinates and real heights. Finally, a geometric transformation converts the network output into real-world 3D lane points.
Liu et al. proposed CLGo [52], a two-stage framework for predicting 3D lanes and camera pose from a single image. The first stage focuses on camera pose estimation and introduces an auxiliary 3D lane task and geometry constraints for multi-task learning. The second stage targets the 3D lane task and uses the previously estimated pose to generate top-view images for accurate 3D lane prediction. PersFormer [12] introduces the first transformer-based approach for 3D lane detection and proposes a novel architecture called Perspective Transformer, as shown in Fig. 4. This transformer-based architecture enables spatial feature transformation, allowing for accurate detection of 3D lane lines. Moreover, the proposed framework has the unique capability to simultaneously handle both 2D and 3D lane detection tasks, providing a unified solution. Additionally, the paper presents OpenLane, a large-scale 3D lane detection dataset built upon the influential Waymo Open dataset [75]. OpenLane is the first dataset of its kind to offer high-quality annotations and diverse real-world scenarios, making it a valuable resource for advancing research in the field.
Figure 4: PersFormer pipeline. The key is to understand the spatial feature transformation from the front view to the BEV space, aiming to enhance the representativeness of the resulting BEV features at the target point by taking into account the local context around the reference point.
In [107], researchers introducesd the largest real-world 3D lane detection dataset, ONCE-3DLanes dataset, and provides a more comprehensive evaluation metric to rekindle interest in this task in real-world scenarios. Additionally, the paper proposes a method called SALAD, which enables direct generation of 3D lane layouts from front view images without the need for converting the feature map into the bird’s-eye view (BEV), the network architecture of SALAD is shown in Fig. 5. This paper [45] proposes a novel loss function that utilizes the geometry structure prior of lanes in 3D space, enabling stable reconstruction from local to global and providing explicit supervision. It introduces a 2D lane feature extraction module that leverages direct supervision from the top view, ensuring maximum retention of lane structural information, especially in distant areas, the overrall pipeline is shown is Fig. 7. Additionally, the paper presents a task-specific data augmentation method for 3D lane detection to address the imbalanced data distribution of ground slope and camera pose, enhancing generalization on rare cases. Bai et al. proposed CurveFormer [7], a Transformer-based algorithm for 3D lane detection. In this paper, researchers formulate the queries in the decoder layers as a dynamic anchor point set, and utilize a curve cross-attention module to compute the similarity between the queries and image features. Besides, they also introduce a context sampling unit to predict offsets by combining reference features and queries, guiding the learning of sampling offsets. Ai et al. proposed WS-3D-Lane [3], which for the first time proposed a weakly supervised 3D lane detection method using only 2D lane labels, and outperformed the previous 3D-LaneNet [24] method in evaluation. In addition, the author proposes a camera pitch self-calibration method, which can calculate the camera pitch angle online in real time, thereby reducing the error caused by the pitch angle change between the camera and the ground plane caused by the uneven road surface. In BEV-LaneDet [89], the author proposed virtual camera, a novel data prediction processing module, to unify the extrinsic parameters of cameras, and the consistency of data distribution, the authors proposed keypoint representation, a simple yet effective representation of 3D lane structures. In addition, the MLP Based Spatial Transformation Pyramid was proposed, a lightweight structure that realizes the conversion from multi-angle visual features to BEV features. Huang et al. proposed the Anchor3DLane [31] framework, which directly defines anchor points in 3D space and directly regresses 3D lanes from the front view without introducing a top view, as shown in Fig. 6. The authors also proposed a multi-frame extension of Anchor3DLane to exploit well-aligned temporal information and further improve performance. Furthermore, a global optimization method is developed to fine-tune the lane lines by utilizing the lane iso-width property.
Figure 5: The architecture of SALAD. The input image is encoded into deep features by the backbone, and then two branches, the semantic awareness branch and the spatial contextual branch, decode the features to extract the spatial information of the lane and create the segmentation mask. Subsequently, the 3D reconstruction process integrates this information to ultimately derive the 3D lane positions in the real-world scene.
Figure 6: The architecture of Anchor3DLane. Utilizing a front-view input image, we employ a CNN backbone and a Transformer layer to extract the visual feature F initially. Subsequently, 3D anchors are projected to sample their features from F based on camera parameters. Following this, a classification head and a regression head are utilized to make the final predictions. The lane predictions can also be used as new 3D anchors for iterative regression.
Figure 7: The framework includes a top-view segmentation network and an anchor-based 3D lane regressor. Features from a front-view image are projected onto the top view to create the segmentation mask under top-view supervision. The anchor representation of 3D lanes is estimated from the top-view lane mask. Additionally, we incorporate the geometry prior from 3D to 2D projection for 3D lane reconstruction.
Li et al. proposed a method [45] to directly extract top view lane information from front-view images, reducing the structural loss in the 2D lane representation. The overall pipeline of this method is shown in the Fig. 7. In this paper, the authors considered 3D lane detection as a reconstruction problem from 2D images to 3D space. They proposed that explicit imposition of the geometric prior of 3D lanes during the training process is essential for fully utilizing the structural constraints of inter-lane and intra-lane relationships, and for extracting the height information of 3D lanes from the 2D lane representation. The authors analyzed the geometric relationship between 3D lanes and their 2D representation, and proposed an auxiliary loss function based on the geometric structure prior. They also demonstrated that explicit geometric supervision can enhance noise elimination, outlier rejection, and structure preservation for 3D lanes. Bai et al. proposed CurveFormer [7] and CurveFormer++ [8], which are single-stage methods based on Transformer that can directly calculate the parameters of 3D lanes and can bypass the challenging view transformation step. Specifically, they formalized 3D lane detection as a curve propagation problem using curve queries. A 3D lane query is represented by a dynamic and ordered set of anchor points. By using queries with curve representation in the Transformer decoder, the 3D lane detection results are iteratively refined. Additionally, they introduced a curve cross-attention module to calculate the similarities between curve queries and image features. Furthermore, a context sampling module is provided to capture more relevant image features of a curve query, further enhancing the performance of 3D lane detection. Similar to [66], Li et al. proposed GroupLane [48], a row-wise classification-based method for 3D lane detection. GroupLane’s design consists of two groups of convolutional heads, each corresponding to a lane prediction. This grouping decouples the interaction of information among different lanes, reducing optimization difficulty. During training, predictions are matched with lane labels using a single-win one-to-one matching (SOM) strategy, which assigns predictions to the most suitable labels for loss computation. To address the challenge posed by the inevitable depth ambiguity in monocular images causing misalignment between the constructed surrogate feature map and the original image for lane detection, Luo et al. proposed a novel LATR model [53]. This is an end-to-end 3D lane detector that uses 3D-aware front-view features without the need for transformed view representation. Specifically, LATR detects 3D lanes through cross-attention based on query and key-value pairs constructed using a lane-aware query generator and dynamic 3D ground positional embedding. On one hand, each query is generated based on 2D lane-aware features and adopts a hybrid embedding to enhance lane information. On the other hand, 3D spatial information is injected as positional embedding from an iteratively updated 3D ground plane. To address the issue of inaccurate view transformation caused by the neglect of road height changes in the process of transforming image-view features into bird’s-eye view using inverse perspective mapping (IPM), Chen et al. proposed an efficient transformer for 3D lane detection [15]. Unlike conventional transformers, the model includes a decomposed cross-attention mechanism that can simultaneously learn lane and bird’s-eye view representations. This approach allows for more accurate view transformation compared to IPM-based methods and is more efficient. Previous studies assumed that all lanes are on a flat ground plane. However, Kim et al. argued that algorithms based on this assumption have difficulty in detecting various lanes in actual driving environments and proposed a new algorithm, D-3DLD [40]. Unlike previous methods, this approach expands rich contextual features from the image domain to 3D space by utilizing depth-aware voxel mapping. Additionally, the method determines 3D lanes based on voxelized features. The authors designed a new lane representation combined with uncertainties and predicted the confidence intervals of 3D lane points using Laplace loss. Li and colleagues proposed a lightweight method [46] that uses MobileNet [29] as the backbone network to reduce the demand for computational resources. The proposed method includes the following three stages. First, the MobileNet model is used to generate multi-scale front-view features from a single RGB image. Then, a perspective transformer calculates the bird’s eye view (BEV) features from the front-view features. Finally, two convolutional neural networks are used to predict the 2D and 3D coordinates and their respective lane types. In paper [26], Han et al. argued that the curve-based lane representation may not be well-suited for many irregular lane lines in real-world scenarios, which can lead to performance gaps compared to indirect representations such as segmentation-based or point-based methods. In this paper, the authors proposed a new lane detection method, which can be decomposed into two parts: curve modeling and ground height regression. Specifically, a parameterized curve is used to represent lanes in the bird’s eye view space to reflect the original distribution of lanes. For the second part, since ground heights are determined by natural factors such as road conditions and are not comprehensive, the ground heights are regressed separately from the curve modeling. Additionally, the authors have designed a new framework and a series of losses to unify the 2D and 3D lane detection tasks, guiding the optimization of models with or without 3D lane labels.
An intuitive summary of these methods is shown in Table II, which includes a description of the methods, the datasets used, the open-source status, and the network architecture.
Figure 8: Several open-source algorithms’ visualization results on the ApolloSim dataset, with 3D-LaneNet, Gen-LaneNet, CLGo, WS-3D-Lane, and Anchor-3D-Lane algorithms displayed from top to bottom. In the images, the red lines represent the ground truth of the lane lines, while the blue lines represent the lane lines predicted by the algorithms.
TABLE I: LIST OF USED VARIABLES
Variable Description
$Accuracy$ The proportion of correctly classified samples to the total number of samples.
$Recall$ The proportion of samples classified as positive and predicted correctly to the total number of true positive samples.
$Precision$ The proportion of samples classified as positive and predicted correctly to the total number of samples classified as positive.
$F-Score$ The harmonic mean of precision and recall.
$\beta$ The parameter used to measure the importance of precision and recall in F-Score.
$AP$ Average Precision.
$CD$ Chamfer distance.
$L^{p}$ The predicted lane.
$L^{g}$ The ground-truth lane.
$P_{gi}$ Equals to $(x_{gi},y_{gi},z_{gi})$, a point of $L^{g}$.
$P_{pj}$ Equals to $(x_{pj},y_{pj},z_{pj})$, a point of $L^{p}$.
$IoU$ Intersection over Union (IoU), used to measure the degree of overlap between predicted bounding boxes and ground truth bounding boxes.
$mIoU$ The average value of IoU.
$MSE$ Mean Square Error.
$MAE$ Mean Absolute Error.
$CE$ Cross-Entropy,a commonly used loss function for evaluating the performance of classification models.
$BCE$ Binary Cross-Entropy, a common loss function used in binary classification tasks.
$y$ Ground truth value.
$\hat{y}$ Network’s predicted value.
$p_{i}$ The probability of the model predicting the positive class.
$p_{t}$ The probability predicted for the actual class, used in Huber loss.
$p_{ic}$ The predicted probability of sample $i$ belonging to class $c$.
$y_{ic}$ The ground true label of sample $i$ for class $c$.
$log$ Logarithmic function.
$\gamma$ TA threshold parameter in Huber loss, defining the point where the loss function transitions from quadratic to linear.
||$\cdot$|| ⋅ | The size of the set (usually the number of pixels).
$\cup$ Union operation.
$\cap$ Intersection operation.
$S$ The ground truth region, used in IoU loss.
$\hat{S}$ The region predicted by the network, used in IoU loss.
$N$ The sample size.
$C$ The number of categories.
III PERFORMANCE EVALUATION OF 3D LANE DETECTION
This section will discuss the performance evaluation of monocular 3D lane detection models. Herein, we explain the evaluation metrics, different types of objective functions, analyze the computational complexity, and finally provide quantitative comparisons of various models. The nomenclature of the used variables is given in Table LABEL:tab:variables.
First, we present the visualization results of 3D lane line detection. As some algorithms are not open-sourced, we have only performed visualization tests on the ApolloSim dataset using a few open-source algorithms. These algorithms have been trained on the ApolloSim dataset, and the visualization results are shown in Fig. 8, where the red lines represent the predicted lane lines, and the blue lines represent the ground truth lane lines.
Next, we will introduce the evaluation metrics, the loss function used for training the algorithms, and the quantitative test results of 3D lane line detection on public datasets.
III-A Evaluation Metrics for 3D Lane Detection
Building only predictive monoculdar 3D lane detection models is not a wise and trustworthy decision for safe AD unless it is tested on unseen data. Most models evaluate their performance on a disjoint set of the same dataset used for training, i.e., the test data remains new to the training model. Deep learning models for monocular 3D lane detection are evaluated using some general metrics to assess the best results based on true values. For monocular 3D lane detection tasks, different types of evaluation metrics are available for these tasks, which we review next as follows:
III-A1 Accuracy
$Accuracy=\frac{N_{TP}+N_{TN}}{N_{TP}+N_{FP}+N_{TN}+N_{FN}}$ (1)
III-A2 Recall
$Recall=\frac{N_{TP}}{N_{TP}+N_{FN}}$ (2)
III-A3 Precision
$Precision=\frac{N_{TP}}{N_{TP}+N_{FP}}$ (3)
III-A4 F-score
$F-score=(1+\beta^{2})\times\frac{Precision*Recall}{\beta^{2}\times Precision+Recall}$ (4)
III-A5 AP
$AP=\sum_{1}^{N}p(k)\Delta r(k)$ (5)
III-A6 X Error
$X_{Error}=\frac{1}{N}\sum_{i=1}^{N}\sqrt{(x_{i}-\hat{x_{i}})^{2}}$ (6)
III-A7 Z Error
$Z_{Error}=\frac{1}{N}\sum_{i=1}^{N}\sqrt{(z_{i}-\hat{z_{i}})^{2}}$ (7)
III-A8 CD
$\begin{cases}CD=\frac{1}{N}\sum_{i=1}^{N}||P_{gi}-\hat{P_{gj}}||_{2},\\ \hat{P_{pj}}=\min_{P_{p_{j}}\in L^{p}}||P_{pj}-P_{gi}||_{2}.\end{cases}$ (8)
III-B Loss Functions for 3D Lane Detection
In the task of monocular 3D lane detection, the common basic loss functions include the following:
III-B1 MSE loss
MSE loss is one of the most commonly used loss functions, which calculates the squared difference between the model’s predicted values and the true values, and then takes the average. Its mathematical expression is:
$L_{MSE}=\frac{1}{N}\sum_{i=1}^{N}(y_{i}-\hat{y}_{i})^{2}$ (9)
III-B2 MAE loss
MAE loss is another commonly used loss function, which calculates the absolute difference between the model’s predicted values and the true values, and then takes the average. Its mathematical expression is:
$L_{MAE}=\frac{1}{N}\sum_{i=1}^{N}|y_{i}-\hat{y}_{i}|$ (10)
III-B3 Huber Loss
Huber Loss combines the advantages of MSE and MAE, making it more robust to outliers. Its mathematical expression is:
$L_{Huber}=\begin{cases}\frac{1}{2}(y-\hat{y})^{2},&\text{if }|y-\hat{y}|\leq\delta\\ \delta(|y-\hat{y}|-\frac{1}{2}\delta),&\text{otherwise}\end{cases}$ (11)
III-B4 Cross-Entropy Loss
Cross-Entropy is commonly used for classification tasks but can also be applied to regression tasks. In lane detection, the problem can be transformed into a classification task by determining whether a pixel belongs to a lane based on its position. Its mathematical expression is:
$L_{CE}=-\frac{1}{N}\sum_{i=1}^{N}\sum_{c=1}^{C}y_{ic}\log(p_{ic})$ (12)
III-B5 Binary Cross-Entropy Loss
Binary Cross-Entropy loss is commonly used for training binary classification tasks, aiming to minimize the loss function to improve the accuracy of the model’s predictions on binary classification samples. It is widely applied in deep learning for tasks such as image classification, text classification, and more.. Its mathematical expression is:
$L_{BCE}=-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\log(p_{i})+(1-y_{i})\log(1-p_{i})\right]$ (13)
III-B6 Focal Loss
Focal Loss is a loss function designed to address the class imbalance problem commonly encountered in tasks such as object detection or semantic segmentation, where the number of examples from one class greatly outnumbers the number of examples from another. This class imbalance can lead to models being biased towards the majority class, resulting in poor performance, especially on the minority class.
$L_{Focal}=-(1-p_{t})^{\gamma}\log(p_{t})$ (14)
III-B7 IoU Loss
IoU Loss is based on Intersection over Union (IoU) and is used in object detection and segmentation tasks to measure the overlap between the predicted regions by the model and the ground truth regions.
$L_{IoU}=1-\frac{|S\cup\hat{S}|}{|S\cap\hat{S}|}$ (15)
Different methods utilize specific loss functions differently, but essentially, most are variations or combinations of the basic loss functions mentioned above. Additionally, the Hungarian algorithm [100] is often used to match predicted lanes with ground truth lanes.
III-C Quantitative Analysis of Monocular 3D Lane Detection Models
This section elaborates on the quantitative empirical analysis of monocular 3D lane detection methods surveyed in this paper, which contribute to the automation of autonomous driving. For quantitative assessment, we utilize four evaluation metrics to examine the performance of each monocular 3D lane detection method on the ApolloSim dataset: AP, F-Score, x error and z error, and we report the results in Table IV. On the Openlane dataset, we evaluated the F-Score of each model, as shown in Table V. On the ONCE-3DLane dataset, we evaluate four metrics, namely: F-Score, Precision, Recall, and CD error, the results are reported om Table VI.
Additionally, computational efficiency is taken into account by reporting the Frames Per Second (FPS) achievable by each method during the inference process. The overall runtime of these models is reported in Table IV,V and VI. In some cases, the time complexity indicators of the models are reported from their methods, while in other cases, we ran the reported models from their publicly available repositories using our experimental resources. The system’s CPU configuration includes an Intel(R) Core i9-12900K CPU processor running on the Ubuntu 20.04 operating system, while the GPU used in the experiments is an NVIDIA GeForce RTX 3080Ti GPU with 12 GB of graphics memory, In the table, we specified the hardware used for inference for each method.
IV DATASETS
In vision tasks based on deep learning, another equally important component is the dataset. In this section, we will introduce the datasets currently available for monocular 3D lane line detection tasks. Some of these datasets are open-sourced and popular, frequently used by the community, while others are only described in papers and not open-sourced. We have compiled an intuitive table of all existing monocular 3D lane line detection datasets, as shown in Table III.
IV-A Apollo 3D Lane Synthetic Dataset
The Apollo 3D Lane Synthetic Dataset [25] is a robust synthetic dataset comprising 10,500 frames of high-resolution 1080 × 1920 monocular RGB images, built using the Unity 3D engine. Each frame is accompanied by corresponding 3D lane labels and camera pitch data. The dataset, based on Silicon Valley in the United States, spans a wide array of environments, including highways, urban areas, residential neighborhoods, and downtown settings. The dataset’s images encapsulate a broad range of daytime and weather conditions, various traffic/obstacle scenarios, and different road surface qualities, thereby providing the dataset with a high degree of diversity and realism.
TABLE II: A detailed description of some representative monocular 3D lane detection algorithms, including methodology, open-sourcing status, utilized datasets, and network architecture.
Ref.
Year
Methodology
Dataset
Code
Network Architecture   Need BEV Trans.?
[24]
2018 Combining novel network architecture, IPM, and anchor-based lane representation to achieve accurate 3D lane detection. 3D-Lanes, Synthetic-3D-Lanes
   N/A
   CNN                 ✔
[25]
2020 Using a two-stage framework that decouples the learning of the image segmentation subnetwork and the geometry encoding subnetwork, reducing the amount of 3D lane labels required for training.
Apollo Synthetic.
   link
   CNN                 ✔
[17]
2020 Introducing uncertainty estimation, utilizing a semi-local, Bird’s Eye View (BEV), tile representation to break down lanes into simple segments, and subsequently clustering them into complete lanes.
Synthetic-3D-Lanes, 3D-lanes.
   N/A
   CNN                 ✔
[16]
2020 An anchor-free method, which can detect complex lane topologies such as splits, merges, short lanes, and perpendicular lanes.
Synthetic-3D-Lanes, Real-3D-lanes
   N/A
   CNN                 ✔
[52]
2021 Presenting a two-stage framework, leveraging auxiliary 3D lane tasks and geometry constraints in the first stage, followed by accurate 3D lane prediction in the second stage using estimated pose.
Apollo Synthetic.
   link
   CNN                 ✔
[12]
2022 The paper introduces a novel Transformer-based spatial feature transformation module and presents OpenLane, a large-scale real-world 3D lane dataset.
Apollo Synthetic, OpenLane.
   link
Transformer           ✔
[107]
2022 The paper introduces the ONCE-3DLanes dataset and proposes SALAD, an extrinsic-free, anchor-free method for regressing 3D lane coordinates in image view.
ONCE-3DLanes.
   link
   CNN                 ✘
[33]
2022 The paper proposes Att-Gen-LaneNet, a two-stage network for robust 3D lane detection. It integrates the Efficient Channel Attention (ECA) and Convolutional Block Attention Module (CBAM) to improve segmentation and generalization.
Apollo Synthetic.
   N/A
   CNN                 ✔
[45]
2022 The paper proposes an approach for monocular 3D lane detection by leveraging the geometry structure and explicit supervision based on the structure prior.
Apollo Synthetic.
   N/A
   CNN                 ✔
[46]
2022 The method comprises three stages: generating multiscale front-view features, calculating bird’s eye view (BEV) features, and predicting 2D and 3D lanes.
OpenLane.
   N/A
   CNN                 ✔
[7]
2022 The method formulates 3D lane detection as a curve propagation problem, enhancing performance by leveraging curve queries, a curve cross-attention module, and a context sampling module to capture relevant image features.
Apollo Synthetic, OpenLane.
   N/A
Transformer           ✘
[3]
2022 The paper introduces a method training 3D lanes with only 2D lane labels, leveraging assumptions of constant lane width and equal height on adjacent lanes for indirect supervision of 3D lane heights.
Apollo Synthetic, ONCE-3DLanes.
   N/A
   CNN                 ✔
[90]
2022 The paper introduces a Virtual Camera for spatial relationship consistency, a Key-Points Representation module for diverse 3D lane structures, and a Spatial Transformation Pyramid for front-view to bird’s eye view feature transformation.
Apollo Synthetic, OpenLane.
   link
   CNN                 ✔
[95]
2022 Utilizing multi-frame input and removing the dependency on camera pose, the model integrates a multi-frame pre-alignment layer, spatio-temporal attention module, and 3DLane IOULoss.
OpenLane, ONCE-3DLanes.
   N/A
   CNN                 ✔
[31]
2023 The paper introduces Anchor3DLane, a BEV-free method predicting 3D lanes directly from front-viewed (FV) representations using defined 3D lane anchors.
Apollo Synthetic, OpenLane, ONCE-3DLanes.
   link
   CNN                 ✘
[40]
2023 The paper introduces a method that employs depth-aware voxel mapping, extends contextual features to 3D space, and predicts confidence intervals using Laplace loss.
LLAMAS-3DLD, Pandaset-3DLD, Apollo Synthetic.
   N/A
   CNN                 ✔
[15]
2023 The paper proposes an efficient transformer for 3D lane detection using a decomposed cross-attention mechanism to learn lane and bird’s-eye-view (BEV) representations concurrently.
OpenLane, ONCE-3DLanes.
   N/A
Transformer           ✘
[48]
2023 Similar to Paper [66], this is the first work on 3D lane detection utilizing fully convolutional heads based on row-wise classification.
OpenLane, ONCE-3DLanes, OpenLane-Huawei.
   N/A
   CNN                 ✔
[53]
2023 The paper introduces LATR, an end-to-end 3D lane detection model utilizing 3D-aware front-view features and cross-attention mechanism with lane-aware query generator.
OpenLane, Apollo Synthetic, ONCE-3DLanes.
   link
Transformer           ✘
[26]
2023 The paper introduces a novel lane detection method, dividing it into curve modeling and ground height regression, within a unified framework for both 2D and 3D lane detection.
OpenLane, ONCE-3DLanes, TuSimple, CULane.
   N/A
   CNN                 ✘
[8]
2024 The approach models the 3D detection task as a curve propagation problem, where each lane is represented by a curve query with a dynamic and ordered anchor point set.
OpenLane, ONCE-3DLanes.
   N/A
Transformer           ✘
The dataset is segmented into three distinct scene categories: Balanced Scenes, Rarely Observed Scenes, and Scenes with Visual Variations. Balanced Scenes serve as a comprehensive and unbiased dataset for benchmarking standard driving scenarios. Rarely Observed Scenes are used for testing an algorithm’s adaptability to rarely encountered situations involving complex urban maps with drastic elevation changes and sharp turns. Scenes with Visual Variations are designed to assess how algorithms perform under different illumination conditions by excluding specific daytime periods during training and focusing on them during testing. The camera’s intrinsic parameters in the dataset are fixed, with the camera heights ranging between 1.4 to 1.8 meters and pitch angles spanning from 0 to 10 degrees.
TABLE III: A brief introduction to the datasets in monocular 3D lane line detection, including their source papers, parent dataset, data type, resolution, data size, the year in which they were created, availability, and the link to their open-source repositories.
Dataset Parent Dataset Data Type Resolution Total Images Year Availability Source
ApolloSim [25] - Synthetic (Unity 3D) 1920 × 1080 10.5K 2020 link
OpenLane [12] Waymo [74] Real-world 1920 × 1280 200K 2022 link
ONCE-3DLanes [107] ONCE [61] Real-world 1920 × 1020 211K 2022 link
Synthetic-3D-Lanes [24] - Synthetic (Blender) 360 × 480 306K 2019 -
Real-world 3D-Lanes [24] - Real-world / 85K 2019 -
LLAMAS-3DLD [40] LLAMAS [9] Real-world 1276 × 717 6.6K 2023 -
Pandaset-3DLD [40] Pandaset [104] Real-world 1920 × 1080 6K 2023 -
IV-B OpenLane
OpenLane [12] stands as the first large-scale, real-world 3D lane detection dataset, boasting over 200,000 frames and 880,000 meticulously annotated lanes. Built on the foundation of the influential Waymo Open dataset, OpenLane utilizes the same data format, evaluation pipeline, and a 10Hz sampling rate, courtesy of 64-beam LiDARs, over a 20-second span. This dataset provides exhaustive details for each frame, inclusive of camera intrinsics and extrinsics, along with lane categories, which encompass 14 distinct types such as white dashed lines and road curbsides. Almost 90% of the total lanes are constituted by double yellow solid lanes, and single white solid and dashed lanes. The OpenLane dataset epitomizes a real-world scenario, duly highlighting the long-tail distribution problem. OpenLane includes all lanes in a frame, even accommodating those in the opposite direction, provided there is no dividing curbside. Due to intricate lane topologies like intersections and roundabouts, individual frames can house up to 24 lanes. Approximately 25% of the frames harbor more than six lanes, surpassing the maximum in most current lane datasets. In addition to this, the dataset provides annotations for scene tags such as weather and locations, along with the closest-in-path object (CIPO) - defined as the most pertinent target with respect to the ego vehicle. This ancillary data proves invaluable for subsequent modules in planning and control, and not merely perception. The 3D ground truth of OpenLane is synthesized using LiDAR, thus ensuring high precision and accuracy. The dataset is split between a training set containing 157,000 images and a validation set comprising 39,000 images.
IV-C ONCE-3DLanes
The ONCE-3DLanes dataset [107] manifests as another practical 3D lane detection dataset, meticulously derived from the ONCE autonomous driving repository. This collection encompasses 211,000 visuals captured by a forward-facing camera, complemented by their matching LiDAR point cloud data. Showcasing a wide spectrum of scenes across varied times of day and weather settings, such as sunlit, overcast, and rainy conditions, the dataset casts light on multiple terrains like city centers, residential areas, highways, bridges, and tunnels. This diversity fortifies the dataset as a crucial resource for developing and validating resilient 3D lane detection models under an assortment of real-world scenarios. The dataset is organized into three segments: 3,000 scenes for validation, 8,000 scenes for testing, and the residual 5,000 scenes reserved for training. The training component is supplemented with an extra 200,000 unlabeled scenarios to thoroughly harness the raw data. While the dataset supplies intrinsic camera parameters, it omits extrinsic camera parameters.
IV-D Other Datasets
The paper “3D-LaneNet: End-to-End 3D Multiple Lane Detection” [24] presents two distinct datasets: the Synthetic-3D-Lanes dataset and the 3D-Lanes dataset. The Synthetic-3D-Lanes dataset, created via the open-source graphics engine, Blender, consists of 300K training and 5K testing examples, each featuring a 360×480 pixel image alongside associated ground truth parameters such as 3D lanes, camera height, and pitch. Exhibiting substantial diversity in lane topography, object placement, and scene rendering, this dataset provides a valuable resource for method development and ablation studies. Complementarily, the 3D-Lanes dataset is a real-world ground truth labeled data collection, compiled through the utilization of a multi-sensor setup - a forward-looking camera, a Velodine HDL32 lidar scanner, and a high-precision IMU. This dataset was assembled from six separate drives, each on a different road segment, collectively spanning nearly two hours of driving. With the aid of Lidar and IMU data, aggregated lidar top view images were generated and used alongside a semi-manual annotation tool to establish the ground truth. In total, 85,000 images were annotated, of which 1,000, derived from a separate drive, were designated as the test set, with the remainder serving as the training set. The 3D-Lanes dataset proves instrumental in validating the transferability of the proposed approach to real-world data and enabling qualitative analysis. While the Synthetic-3D-Lanes dataset is made available to the research community, the real-world 3D-Lanes dataset remains proprietary and is not publicly accessible. It’s worth noting that despite its availability, the Synthetic-3D-Lanes dataset has not seen wide adoption for benchmarking in subsequent studies in the field.
V 3D LANE DETECTION IN AD: CHALLENGES AND DIRECTIONS
The datasets introduced above encompass a variety of publicly available road scenes. Current mainstream research primarily focuses on favorable daytime scenes that are suitable for and conducive to three-dimensional lane detection, with sufficient lighting and favorable weather conditions. Many car companies and original equipment manufacturers in the industry possess a large volume of data; however, they are reluctant to share it publicly due to concerns related to intellectual property, industrial competition, and General Data Protection Regulations (GDPR). As a result, the lack of sufficient annotated data for accurately understanding dynamic weather conditions, such as nighttime, smoggy weather, and edge cases, remains a challenging task in autonomous driving research.
This research area is among the challenges that the community has not yet adequately addressed. In this section, we present critical viewpoints on the current state of three-dimensional lane detection in autonomous driving, summarize a series of challenges, and provide research direction recommendations to help the community make further progress and effectively overcome these challenges.
V-A Open Challenges
Although significant research has been conducted by researchers in the field of autonomous driving (AD), and the AD industry is flourishing, there are still open challenges that require researchers’ attention to achieve fully intelligent autonomous driving. These challenges have been individually discussed with support from relevant literature:
V-A1 Coarse-Structured Information
Coarse-Structured Information: Most of the datasets introduced in the literature for 3D lane detection in autonomous driving (AD) are recorded in normal and well-structured infrastructures of advanced cities. The currently developed deep learning models may achieve the best results over structured datasets, but they generalize poorly in many unstructured environments. This aspect of AD demands further attention in terms of data collection, as well as the inclusion of new and effective representation mechanisms in deep learning models.
V-A2 Uncertainty-Aware Decisions
A largely overlooked aspect in lane detection and AD decision-making is the confidence with which models make predictions on input data. However, the confidence of model outputs plays a crucial role in ensuring the safety of autonomous driving. The inherently uncertain nature of the vehicular surroundings seems to have not convinced the community to delve into this matter, as the current methodological trends focus solely on predictive scores. Fortunately, confidence estimation has recently gained attention in the community [5, 63]. Nevertheless, elements from evidential deep learning [71], Bayesian formulations of deep neural networks [38], simpler mechanisms to approximate the output confidence of neural networks (e.g., Monte Carlo dropout [21] or ensembles [44]), and other assorted methods for uncertainty quantification [1] should be progressively incorporated as an additional yet crucial criterion for decision making. When dealing with complex environments, where there is a lack of data that fully represents all possible scenes, the model outputs a significant amount of epistemic uncertainty. Without considering confidence as an additional factor for AD or with current studies focused solely on predictive and/or computational efficiency aspects, there is no guarantee that the emerging 3D lane detection models in the scientific community will be practically useful and transferable to the industry.
V-A3 Weakly supervised Learning Strategies
In the current deep learning-based algorithm models, the majority rely on fully supervised learning strategies, which have high demands for labeled data. This is particularly challenging in the field of 3D lane detection, as general visual sensor data lacks depth information. It is difficult to simply assign 3D information to lanes based on images alone, and alternative sensors such as LiDAR are required to obtain the 3D lane information. This leads to the costly and labor-intensive task of annotating 3D lane data. However, fortunately, both the academic and industrial communities have recognized this issue, and weakly supervised learning strategies have received extensive research and attention in the field of deep learning [reference]. However, in the specific branch of 3D lane detection, there is currently limited research focusing on weakly supervised learning strategies. If we can effectively leverage self-supervised/weakly supervised learning strategies, it would significantly reduce data collection costs and allow for more training data to enhance the performance of lane detection algorithms, thereby further promoting the advancement of the autonomous driving industry.
V-B Future Directions
V-B1 Video-Based 3D Lane Detection for Autonomous Driving
Drawing from the advancements in video-based object detection [102], semantic segmentation [84, 60, 30], and 2D lane detection [34, 112, 111], it’s evident that incorporating video-based techniques significantly improves the precision and dependability of 3D lane detection systems. The core advantage of video-based methods is their capacity to harness temporal data, providing a dynamic view that static images lack. This dynamic perspective is particularly vital in comprehending and predicting complex driving situations in three-dimensional spaces, where the intricacies of lane placement and vehicular interactions intensify. Methods like Recursive Video Lane Detection (RVLD) [34] demonstrate the power of video in capturing ongoing lane shifts and variations over time, a feature immensely beneficial for 3D modeling accuracy. Additionally, the assimilation of video data into these systems allows for an enhanced understanding of spatial dynamics in driving environments, a crucial factor in 3D lane detection where accurate identification of lane positioning and curvature in relation to other road elements is critical. By incorporating sophisticated deep learning techniques used in video-based object detection and semantic segmentation, future iterations of 3D lane detection systems could attain advanced spatial awareness, significantly enhancing the navigation capabilities and safety of autonomous vehicles.
V-B2 Hybrid Approaches and Multimodality
The progression of multi-modal 3D lane detection technologies for autonomous vehicles is greatly accelerated by the incorporation of various sensory inputs, such as cameras, LiDAR, and radar. This integration marks a promising avenue for overcoming the challenges faced by existing camera-reliant systems. This approach, underscored by the successes in multi-modal 3D object detection and semantic segmentation [19], leverages the complementary strengths of each sensor type to enhance detection accuracy and reliability, particularly in challenging environmental and complex driving scenarios [13, 49]. Reflecting on pioneering models such as “Deep Multi-Sensor Lane Detection” [6] and “M^2-3DLaneNet” [54], which have effectively utilized multi-sensor inputs to refine lane boundary estimations and demonstrated robust performance against occlusions and variable lighting conditions, the potential for significant advancements in lane detection technology is evident. The future trajectory in this field should emphasize the exploration of advanced data fusion methods, meticulous sensor calibration, and synchronization techniques, alongside the exploitation of emerging technologies like edge computing for real-time multi-modal data processing.
V-B3 Active Learning and Incremental Learning
Active learning [69] in machine learning refers to the ability of a model to adapt and learn over time and encountering new data during the testing phase and after deployment. In real-world environments, vehicles may encounter randomly occurring unfamiliar scenes and lane topology, which may require AI models to make decisions for further actions such as braking or accelerating to achieve sensible driving operations. Therefore, lane detection techniques should allow for an interactive approach to handle various types of scenes and lane topologies, involving human annotators to label unlabeled data instances and human involvement in the training process [39]. There are different types of active learning techniques, such as membership query synthesis [37], where synthetic data is generated and the parameters of the synthetic data can be adjusted based on the structure of the data [2], which originates from the underlying species of the dataset. On the other hand, the ability of 3D lane detection models to incrementally update their captured knowledge with new data is crucial for their sustainability and continuous improvement. We anticipate that both these capabilities of 3D lane detection models in road understanding will become increasingly important in future research.
V-B4 Adverse Weather Conditions
The progress in developing camera-based 3D lane detection systems for autonomous driving is notably hindered by adverse weather conditions, which significantly impair visibility. Events like heavy rainfall, fog, snow, and dust storms can severely impact the functionality of these systems [113]. The core issue stems from the compromised quality of visual data necessary for the precise detection and segmentation of lane markings, resulting in a decline in reliability and an increase in the likelihood of false negatives or positives. This diminishment in the efficacy of such systems not only elevates safety hazards but also constrains the operational scope of autonomous vehicles. However, recent breakthroughs in object detection and segmentation, as illustrated in research works such as ”ACDC: The Adverse Conditions Dataset with Correspondences for Semantic Driving Scene Understanding” [70] and ”Vehicle Detection and Tracking in Adverse Weather Using a Deep Learning Framework” [27], reveal avenues for enhancing 3D lane detection amidst challenging weather conditions. These studies put forth the advantages of employing deep learning algorithms trained on datasets that include a wide array of adverse weather instances, showcasing the significance of effective data augmentation, domain adaptation tailored to specific conditions, and the employment of semantic segmentation techniques. Employing these approaches, the capabilities of camera-based detection systems can be substantially improved to interpret lane markings accurately and ensure safe navigation even under poor visibility, laying down an optimistic pathway for ongoing research and development in the field of autonomous driving technology.
Figure 9: The test results of the current open-source state-of-the-art LATR [53] algorithm under extreme weather conditions or photography scenarios are depicted in the figure above, with red lines indicating the algorithm’s predicted lane lines and blue lines representing the ground truth lane lines. Figures a and d show scenes with strong exposure during the day and at night, respectively; figures b and e depict rainy scenes during the day and night; and figures c and f show foggy scenes during the day and night. From these test scenarios under harsh conditions, it is evident that the current algorithm performs poorly in detecting lane lines, indicating that there is substantial room for improvement.Additionally, there is a lack of datasets for training and testing under such severe conditions.
V-B5 3D Lane Detection with Large Language Model (LMM)
The advent of large language models (LLMs) such as ChatGPT [103] has transformed the field of artificial general intelligence (AGI), demonstrating their remarkable zero-shot abilities in tackling a variety of natural language processing (NLP) tasks using customized user prompts or language instructions.
Computer vision encompasses a set of unique challenges and concepts that stand apart from those found in NLP. Vision foundation models typically adhere to a pre-training and subsequent fine-tuning process [92, 14, 73, 93, 18, 80], which, while effective, entails significant additional costs for adaptation to a range of downstream applications. As depicted in Figure 1a, techniques such as multi-task unification [68, 87, 4, 85] aim to endow systems with a broad set of capabilities, yet they often fail to transcend the constraints of predetermined tasks, leaving a noticeable capability deficit in open-ended tasks when juxtaposed with LLMs. The advent of visual prompt tuning [32, 108, 110, 94] provides a novel method for delineating specific vision tasks, like object detection, instance segmentation, and pose estimation, through the use of visual masking. Nonetheless, there is currently no work that combines LLMs with 3D lane line detection. As large language models become more widespread and their capabilities continue to improve, research into LLM-based lane line detection presents an interesting and promising avenue for future exploration.
Scene Method FPS Device AP(%) F1(%) x err/N(m) x err/F(m) z err/N(m) z err/F(m)
Balanced Scene 3DLaneNet [24] 53 RTX 2080 89.3 86.4 0.068 0.477 0.015 0.202
Gen-LaneNet [25] 60 RTX 2080 90.1 88.1 0.061 0.496 0.012 0.214
CLGo [52] 75 RTX 2080 94.2 91.9 0.061 0.361 0.029 0.250
PersFormer [12] 21 Tesla V100 - 92.9 0.054 0.356 0.010 0.234
Reconstruct from Top [45] 82 GTX 1080Ti 93.8 91.9 0.049 0.387 0.008 0.213
CurveFormer [7] - - 97.3 95.8 0.078 0.326 0.018 0.219
WS-3D-Lane [3] 26.7 RTX 3080Ti 95.7 93.5 0.027 0.321 0.006 0.215
Anchor3DLane [31] 87.3 RTX 2080Ti 97.2 95.6 0.052 0.306 0.015 0.223
BEV-LaneDet [90] 102 Tesla V100 - 98.7 0.016 0.242 0.020 0.216
LATR [53] 17.75 Tesla A100 97.9 96.8 0.022 0.253 0.007 0.202
D-3DLD [40] 22.6 Tesla A100 95.5 93.3 0.039 0.412 0.011 0.234
Rarely Observed 3DLaneNet [24] 53 RTX 2080 74.6 72.0 0.166 0.855 0.039 0.521
Gen-LaneNet [25] 60 RTX 2080 79.0 78.0 0.139 0.903 0.030 0.539
CLGo [52] 75 RTX 2080 88.3 86.1 0.147 0.735 0.071 0.609
PersFormer [12] 21 Tesla V100 - 87.5 0.107 0.782 0.024 0.602
Reconstruct from Top [45] 82 GTX 1080Ti 85.2 83.7 0.126 0.903 0.023 0.625
CurveFormer [7] - - 97.1 95.6 0.182 0.737 0.039 0.561
WS-3D-Lane [3] 26.7 RTX 3080Ti 88.0 86.0 0.070 0.741 0.015 0.585
Anchor3DLane [31] 87.3 RTX 2080Ti 96.9 94.4 0.094 0.693 0.027 0.579
BEV-LaneDet [90] 102 Tesla V100 - 99.1 0.031 0.594 0.040 0.556
LATR [53] 17.75 Tesla A100 97.3 96.1 0.050 0.600 0.015 0.532
Visual Variations 3D-LaneNet [24] 53 RTX 2080 74.9 72.5 0.115 0.601 0.032 0.230
Gen-LaneNet [25] 60 RTX 2080 87.2 85.3 0.074 0.538 0.015 0.232
CLGo [52] 75 RTX 2080 89.2 87.3 0.084 0.464 0.045 0.312
PersFormer [12] 21 Tesla V100 - 89.6 0.074 0.430 0.015 0.266
Reconstruct from Top [45] 82 GTX 1080Ti 92.1 89.9 0.060 0.446 0.011 0.235
CurveFormer [7] - - 93.0 90.8 0.125 0.410 0.028 0.254
WS-3D-Lane [3] 26.7 RTX 3080Ti 95.8 93.6 0.025 0.338 0.005 0.237
Anchor3DLane [31] 87.3 RTX 2080Ti 93.6 91.4 0.068 0.367 0.020 0.232
BEV-LaneDet [90] 102 Tesla V100 - 96.9 0.027 0.320 0.031 0.256
LATR [53] 17.75 Tesla A100 96.6 95.1 0.045 0.315 0.016 0.228
TABLE IV: An in-depth comparison of various methods on the ApolloSim dataset under three different split settings. “N” and “F” are short for near and far respectively.
Method FPS Device All Up & Down Curve Extreme Weather Night Intersection Merge & Split
3D-LaneNet [24] 53 RTX 2080 44.1 40.8 46.5 47.5 41.5 32.1 41.7
GenLaneNet [25] 50 RTX 2080 32.3 25.4 33.5 28.1 18.7 21.4 31.0
PersFormer [12] 21 Tesla V100 50.5 42.4 55.6 48.6 46.6 40.0 50.7
CurveFormer [7] - - 50.5 45.2 56.6 49.7 49.1 42.9 45.4
Anchor3DLane [31] 87.3 RTX 2080Ti 54.3 47.2 58.0 52.7 48.7 45.8 51.7
BEV-LaneDet [90] 102 Tesla V100 58.4 48.7 63.1 53.4 53.4 50.3 53.7
Efficient Transformer [15] 72 Tesla V100 63.8 57.6 73.2 57.3 59.7 57.0 63.9
GroupLane [48] 38.2 RTX 2080 64.1 - - - - - -
LATR [53] 17.75 Tesla A100 61.9 55.2 68.2 57.1 55.4 52.3 61.5
DecoupleLane [26] - - 51.2 43.5 57.3 - 48.9 43.5 -
CurveFormer++ [8] 11.4 RTX 3090 52.7 48.3 59.4 50.6 48.4 45.0 48.1
TABLE V: Comparison of various 3D methods on the OpenLane benchmark, evaluating the F-Score on both the entire validation set and separate sets corresponding to five different scenarios.
Method FPS Device F1(%) Precision(%) Recall(%) CD Error(m)
3D-LaneNet [24] 53 RTX 2080 44.73 61.46 35.16 0.127
Gen-LaneNet [25] 60 RTX 2080 45.59 63.95 35.42 0.121
SALAD [107] - - 64.07 75.90 55.42 0.098
PersFormer [12] 21 Tesla V100 74.33 80.30 69.18 0.074
WS-3D-Lane [3] 26.7 RTX 3080Ti 77.02 84.51 70.75 0.058
Anchor3DLane [31] 87.3 RTX 2080Ti 74.87 80.85 69.71 0.060
Efficient Transformer [15] 72 Tesla V100 80.84 84.50 77.48 0.056
GroupLane [48] 38.2 RTX 2080 80.73 82.56 78.90 0.053
LATR [53] 17.75 Tesla A100 80.59 86.12 75.73 0.052
DecoupleLane [26] - - 75.07 81.19 69.26 0.062
CurveFormer++ [8] 11.4 RTX 3090 77.85 81.06 74.89 0.084
TABLE VI: The performance of 3D lane detection on the ONCE-3DLane dataset. CD error represents the Chamfer distance between the predicted lane line and the ground truth.
V-B6 Towards More Accurate and Efficient 3D Lane Detection Methods for Autonomous Driving
The qualitative performance of current 3D lane detection techniques is shown in Table IV. We can observe that only a few methods manage to strike a balance between model accuracy and inference latency. The experimental results of these methods indicate the need for further improvements to alleviate computational burdens while maintaining their unparalleled performance. Additionally, We added noise to the dataset to simulate adverse weather conditions, and tested the performance of the 3D lane line detection algorithm on these challenging datasets, the performance of the algorithm was not satisfactory, as depicted in Figure 5. We found that these models should also be evaluated for knowledge transferability and generalization across different datasets. Moreover, the reported time complexities in Table [?] suggest that some methods can achieve real-time execution when deployed on GPU devices. Nevertheless, considering the constrained computational resources in today’s autonomous driving telemetry, the focus of 3D lane detection methods should also shift towards computational complexity.
V-B7 3D Lane Detection based on Event Cameras
RGB cameras are limited by their imaging principles, resulting in poor image quality in high-speed or low-light scenes. Fortunately, event cameras can overcome this limitation. Event cameras are visual sensors with high temporal resolution, high dynamic range, low latency, and low energy consumption [99]. Unlike traditional cameras that capture images based on the intensity and color of light, event cameras capture images based on changes in light intensity [22, 23]. Therefore, event cameras can capture images in low-light scenes as long as there are changes in light intensity. As a result, there is significant and extensive research potential in the field of 3D lane detection based on event cameras, including the development of datasets for 3D lane detection with event cameras and Algorithms specifically designed for 3D lane detection using event cameras or a combination of event and RGB cameras.
V-B8 Uncertainty-aware 3D Lane Detection
Over the past few years, DNNs have garnered significant success in numerous computer vision assignments, solidifying their status as indispensable instruments for proficient automated perception. Despite consistently delivering exceptional results across different benchmarks and tasks, there remain several significant hurdles to overcome before widespread implementation. Among the most common and well-known critiques of DNNs is their susceptibility to unreliable performance when confronted with shifts in data distribution levels, highlighting the urgent need to rectify this limitation [20]. Currently, the majority of deep learning models provide deterministic outputs, giving a single result. However, in real-world driving scenarios, it is desirable for models to provide uncertainty estimates for their predictions [62, 38, 57]. Downstream decision-making modules can then use this uncertainty information to make more reasonable and safe driving commands. For example, in the context of 3D lane detection, if the model outputs a lane position with high uncertainty, we should approach the model’s detection result with skepticism, and a conservative driving style should be adopted. Conversely, if the model’s output has low uncertainty, we can have confidence in the algorithm’s prediction and make more confident driving decisions.
VI CONCLUSION
Vision sensors are a critical component of autonomous vehicles, playing a key role in the decision-making process. As one of the fastest-growing fields in recent years, computer vision techniques are used to analyze data captured by vision sensors to obtain useful information such as traffic light detection, traffic sign recognition, drivable area detection, and 3D obstacle perception. With advancements in sensor technology, algorithmic capabilities, and computational power, the use of vision sensor data for autonomous vehicle perception is gaining increasing attention. Monocular-based 3D lane detection, for example, uses a single camera image to obtain the position of lane lines in the 3D physical world, incorporating depth information. Knowing the depth information of lane lines is crucial for safe and comfortable decision-making and planning in autonomous vehicles. While 3D lane information can be obtained using other sensors such as LiDAR, vision sensors play a crucial role in the field of autonomous driving due to their cost-effectiveness and rich structured color information.
3D lane detection based on monocular image has been developed for many years in the field of autonomous driving. However, there is a lack of comprehensive and summarized analysis in the existing literature. In this survey, we reviewed existing lane detection methods, introduced existing three-dimensional lane detection datasets, and discussed the performance comparison of existing lane detection methods on public datasets. We also analyzed the challenges and limitations faced by current three-dimensional lane detection. Our main conclusion is that the research in the field of three-dimensional lane detection based on monocular images has not reached perfection, and current methods have many limitations, which we discussed in detail in the survey and provided relevant suggestions and outlooks. We covered baseline works dealing with deep learning models, their hierarchy for three-dimensional lane detection tasks, and the challenges associated with each model category. Furthermore, we delved into performance evaluation strategies, loss functions, and widely used datasets in the field of autonomous driving for three-dimensional lane detection models. We summarized the work by presenting open challenges and future research directions, along with stimulating baseline references from recent literature.
Lastly, it is undeniable that experts from the intelligent transportation systems community are continuously striving to improve strategies for three-dimensional lane detection to effectively utilize data from vision sensors. Mainstream research is dedicated to enhancing model accuracy through the capabilities of neural networks or exploring novel neural network architectures. However, addressing other challenges is necessary to achieve reliable, trustworthy, and safe autonomous driving. From the perspective of three-dimensional lane detection, these challenges demand more robust models with the ability to anticipate lane occlusions, process coarse-structure information, and provide risk alerts. Additionally, current three-dimensional lane detection models primarily rely on supervised learning, which requires high-quality labeled data. However, annotating three-dimensional lane data is a laborious and time-consuming task. Exploring valuable and challenging approaches, such as self-supervised or weakly supervised learning, to achieve three-dimensional lane detection is an open opportunity for further development in this field. If these opportunities can be fully leveraged in a timely manner, it can drive research in intelligent transportation systems and elevate three-dimensional lane detection to a new level. This would enable unmanned vehicles to be deployed more effectively in real-world environments and support safer, more reliable, and more comfortable travel and logistics services.